# Text-Trained LLMs Can Zero-Shot Extrapolate PDE Dynamics

### v1

@@ A.1.1 Quantization and Reconstruction Implementation
This appendix provides the implementation details for the linear quantization and reconstruction steps described in Section 3. Let $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}}$ denote the floating-point PDE solution evaluated on a uniform spatiotemporal grid with $N_{\mathrm{X}}$ interior spatial points $\{x_{i}\}_{i=1}^{N_{\mathrm{X}}}$ and $N_{\mathrm{T}}+1$ discrete time steps $\{t_{j}\}_{j=0}^{N_{\mathrm{T}}}$. Define
@@
Quantization. Each entry $u(x_{i},t_{j})$ is mapped to
$Q_{i,j}=\begin{cases}500,&u_{\max}=u_{\min},\\ \displaystyle\mathrm{round}\Bigl{(}150+(u(x_{i},t_{j})-u_{\min})\,\frac{850-150}{u_{\max}-u_{\min}}\Bigr{)},&\text{otherwise},\end{cases}$
@@
Reconstruction. To recover an approximate floating-point value $\tilde{u}(x_{i},t_{j})$ from $Q_{i,j}$, we apply a linear reconstruction map that approximates the original values:
$\tilde{u}(x_{i},t_{j})=\begin{cases}u_{\min},&u_{\max}=u_{\min},\\ \displaystyle u_{\min}+(Q_{i,j}-150)\,\frac{u_{\max}-u_{\min}}{850-150},&\text{otherwise}.\end{cases}$
Quantization Error Floor. We report two per-time-step error metrics for the reconstructed solution:
@@
A.1.2 Spline-Based Random Initial Condition Construction
In Section 4, we construct each random initial condition $u_{0}(x)$ by sampling independent values on a fixed grid and then fitting a $C^{2}$ interpolant via cubic splines, yielding a function that is $C^{2}$ on [L,L][-L,L]. This analytic procedure is computationally cheap and helps ensure that any variation in LLM prediction error across different spatial discretizations arises purely from discretization, not from changes in the underlying random sample (see Section 4.1).
Coarse Grid and Sampling. Let the 1D domain be [L,L][-L,L], with Dirichlet boundary values
@@ A.1.2 Spline-Based Random Initial Condition Construction
Spline Interpolant. Use SciPy’s CubicSpline with default “not-a-knot” end conditions to fit
@@
Resampling at a New Grid Resolution. To evaluate different spatial discretizations, choose a new cardinality of interior points $N_{\mathrm{X}}^{\mathrm{new}}$ and set
@@
Parameter Choices. All experiments in Section 4 are conducted with $N_{\mathrm{X}}=14,u_{\mathrm{BC}}=-1,a=-0.5,b=0.5$.
A.2 Evaluation Metrics and Reference Solution Setup
@@
1:Fixed Quantities: Initial conditions $\{\tilde{u}_{0,m}\}_{m=1}^{M}$ and corresponding reference solutions $\{\tilde{u}_{m}(x,t)\}_{m=1}^{M}$ precomputed on a suitably highly refined finite-difference grid using FTCS.
2:Monte Carlo Trials: For $m=1,2,\ldots,M$, run $\mathrm{A}\in\{\mathrm{LLM},\mathrm{Classical~Solver}\}$ given $\{\tilde{u}_{m}(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}-1}$ and obtain prediction $\{\hat{u}_{m}^{\mathrm{A}}(x_{i},t_{N_{\mathrm{T}}})\}_{i=1}^{N_{\mathrm{X}}}$.
@@
$\hskip-28.45274pt\bar{H}^{\mathrm{LLM}}_{m}=-\frac{1}{N_{\mathrm{X}}}\sum_{i=1}^{N_{\mathrm{X}}}\sum_{y\in\mathcal{V}}p_{m}(y\mid x_{i},N_{\mathrm{T}})\log p_{m}(y\mid x_{i},N_{\mathrm{T}}),\qquad\text{(Mean Entropy)}$
@@
In this appendix, we detail the Monte Carlo procedure used to compute the evaluation metrics in Section 4. For completeness, we also report results using the Maximum Absolute Error (MaxAE), which captures the worst-case deviation:
@@
As shown in Figure 6, MaxAE exhibits qualitatively similar trends to the RMSE reported in the error analysis of Section 4. Algorithm 1 summarizes the Monte Carlo procedure used for computing evaluation metrics in the one-step prediction task. The same procedure extends naturally to the multi-step setting, where the metrics are computed identically at each predicted time slice.
@@
Figure 6: Prediction accuracy of Llama-3 models evaluated using the MaxAE metric, under the same experimental setup as in Sections 4.1 and 4.2.
A.3 Results on Additional PDEs
A.3.1 Results on Fisher–KPP, Heat, and Wave Equations
We extend the analysis from Section 4 to three additional PDEs: the Fisher–KPP equation, the heat equation, and the wave equation. All experiments—including one-step prediction, multi-step rollout, and entropy-based uncertainty quantification—are performed under the same setup as in Section 4, [^5] For the Fisher–KPP equation, we set $a=0.2,b=0.8$ (instead of a=0.5,b=0.5a=-0.5,b=0.5), following Appendix A.1.2 so that the initial condition $u(x,0)$ lies within $[0,1]$, consistent with interpreting $u(x,t)$ as a population density. with homogeneous Dirichlet boundary conditions unless noted otherwise. For the wave equation, which is second-order in time, we additionally impose zero initial velocity, i.e., $\partial_{t}u(x,0)=0$, so that the dynamics are fully determined by the initial condition $u(x,0)$. Explicitly, the governing equations are:
@@ A.3.1 Results on Fisher–KPP, Heat, and Wave Equations
For the Fisher–KPP equation, we adopt commonly used parameter values with diffusion coefficient $D=0.002$ and reaction rate $r=1$ (Hasnain & Saqib, 2017; Needham et al., 2025). In the representative results (Figures 710), we present results with thermal diffusivity $k=0.01$ and wave speed $c=0.2$. In Figure 11, we further confirm that the qualitative scaling trends persist across a range of $k$ and $c$ values. For each PDE, we additionally include representative numerical benchmarks to contextualize LLM predictions: FTCS and IMEX for the Fisher–KPP equation, FTCS and BTCS (backward time, centered space) for the Heat equation, and Leapfrog and Crank–Nicolson for the Wave equation (LeVeque, 2007). Beyond Dirichlet boundaries, we also study the heat equation under homogeneous Neumann boundary conditions, where total thermal energy conservation is the key structural property. Notably, LLM rollouts preserve this conservation law, suggesting that ICL can capture deeper invariants of PDE dynamics. Full details are provided in Appendix A.3.2.
As shown in Figures 7, 8, 9, and 10, the qualitative trends closely mirror those observed for the Allen–Cahn equation presented in Section 4. In the one-step prediction setting, accuracy improves systematically with longer temporal context while degrading at finer spatial discretizations. In the multi-step rollout setting, errors accumulate algebraically with the rollout horizon, resembling the global error growth of classical numerical solvers. Entropy-based analysis reveals a consistent three-stage progression in ICL behavior, and prediction uncertainty increases with longer spatial outputs.
@@
Figure 7: One-step prediction error for the Fisher–KPP, Heat, and Wave equations as a function of input context length, under the experimental setup of Section 4.1.
@@
Figure 8: One-step prediction error for the Fisher–KPP, Heat, and Wave equations as a function of output length, under the experimental setup of Section 4.1.
@@
Figure 9: Multi-step prediction error for the Fisher–KPP, Heat, and Wave equations as a function of prediction steps, under the experimental setup of Section 4.2.
@@
Figure 10: Uncertainty analysis of the Fisher–KPP, Heat, and Wave equations, under the experimental setup of Section 4.3.
@@
Figure 11: One-step prediction error with the Llama-3.2-3B model under varying PDE parameters. Shaded regions show 95% confidence intervals over 50 random initial conditions, consistent with the setup in Section 4.
Figure 12: Relative energy deviation $\Delta\mathrm{E}$ over prediction steps for LLM rollouts compared against classical finite difference solvers (FTCS, BTCS) under homogeneous Neumann boundary conditions for the heat equation. Shaded regions denote 95% confidence intervals over 20 random initial conditions. The grid-induced reference is computed by evaluating the total thermal energy with a high-resolution solver and then restricting it to the same coarse spatial grid used in the experimental setup; this represents the minimal deviation expected from discretization alone. Llama-3.1-8B predictions remain close to this reference and exhibit substantially lower conservation error than coarse-grid finite difference solvers across the rollout horizon.
A.3.2 Conservation Properties in the Heat Equation with Neumann Boundaries
Motivation. Beyond continuing spatiotemporal trajectories, an important question is whether LLMs internalize deeper invariants of PDE dynamics. The heat equation with homogeneous Neumann boundary conditions and no internal source term offers a natural test case: it models an insulated rod, where no heat can flow across the boundaries. In this setting, the total thermal energy is conserved for all time (Haberman, 2013). Remarkably, we find that LLM rollouts respect this conservation law more faithfully than coarse-grid finite difference solvers under the same setup, suggesting that ICL captures structural properties of the dynamics rather than performing naive extrapolation.
Conservation Law. The governing PDE is:
@@ A.3.2 Conservation Properties in the Heat Equation with Neumann Boundaries
Relative Energy Deviation. To evaluate conservation in rollouts, we approximate $\mathrm{E}(t)$ using the trapezoidal rule on the spatial grid $\{x_{i}\}_{i=0}^{N_{\mathrm{X}}+1}$. The grid follows the uniform setup in the main section:
@@
Figure 12 reports results under the same setup as the Neumann-boundary heat equation experiments shown in Figure 9. The only modification is that initial conditions are drawn from $a=0,b=1$ (instead of a=0.5,b=0.5a=-0.5,b=0.5), following Appendix A.1.2, so that the conserved energy $\mathrm{E}(0)$ is bounded away from zero. This ensures numerical stability when evaluating relative deviations. We also verified that one-step, multi-step, and uncertainty-evolution analyses under Neumann boundaries with $a=0,b=1$ show consistent qualitative behavior with the results in Appendix A.3.1 (which use a=0.5,b=0.5a=-0.5,b=0.5); to avoid redundancy, we do not reproduce these plots here, but the full results are available in the accompanying GitHub repository. Under this setup, LLM rollouts maintain thermal energy close to its conserved value across time, and notably more faithfully than coarse finite difference solvers operating at the same resolution.
@@
A.4 Results for Instruction-Tuned Llama-3 Variants
@@
Figure 13: Prediction accuracy of instruction-tuned Llama-3 models, using the same experimental setup described in Sections 4.1 and 4.2.
@@
Figure 14: Uncertainty analysis of instruction-tuned Llama-3 models, using the same experimental setup described in Section 4.3 and Figure 5. Top: Mean spatial entropy $\bar{H}$ as a function of context length $N_{\mathrm{T}}$ (left) and output length $N_{\mathrm{X}}$ (right). Bottom: Token-level softmax distributions at the syntax-only stage for Llama-3.1-8B-Instruct. The overall entropy behavior remains consistent with the pretrained base model, with one notable difference: at very short contexts ($N_{\textrm{T}}=2$), the instruction-tuned 8B variant exhibits higher average spatial entropy, reflecting greater uncertainty under minimal context. However, separator tokens (e.g., commas) are still predicted with near-perfect confidence, and this increased uncertainty arises from spatial value tokens more frequently acting as generic placeholders rather than producing deterministically incorrect outputs, as observed in the base model.
@@
We replicate the full experimental setup from Section 4 using instruction-tuned variants of Llama-3 models: Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, and Llama-3.2-1B-Instruct. These models are instruction-tuned for assistant-like chat, whereas the pretrained base models are designed to support a broader range of natural language generation tasks (Grattafiori et al., 2024). All evaluations—one-step prediction, multi-step rollout, and entropy-based uncertainty quantification—are conducted using the same setup as described in Section 4.
@@
As shown in Figure 13, the qualitative trends closely mirror those of the base models. In particular, in one-step prediction settings, accuracy improves with longer input context and degrades with increasing output length. In multi-step prediction settings, predictions exhibit algebraic error accumulation. Similarly, Figure 14 shows that prediction uncertainty undergoes stage-wise transitions as input context increases and grows with output length. Notably, under the same accuracy evaluation setup, the smaller instruction-tuned models, Llama-3.2-3B-Instruct and Llama-3.2-1B-Instruct, show reduced prediction accuracy compared to their pretrained base counterparts. This observation is consistent with prior findings from Gruver et al. (2023), which suggest that alignment procedures such as instruction tuning and Reinforcement Learning with Human Feedback (RLHF) can adversely affect time-series forecasting performance in Llama-2 models. In contrast, we do not observe such a negative impact on the 8B instruction-tuned variant, indicating that newer, larger models may be more robust to the effects of alignment in the context of continuing the spatiotemporal dynamics of PDEs.
A.5 Results for Other Model Families

### v2

@@ 1.1 Quantization and Reconstruction Implementation
This section provides the implementation details for the linear quantization and reconstruction steps described in the Methods section of the main text. Let $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}}$ denote the floating-point PDE solution evaluated on a uniform spatiotemporal grid with $N_{\mathrm{X}}$ interior spatial points $\{x_{i}\}_{i=1}^{N_{\mathrm{X}}}$ and $N_{\mathrm{T}}+1$ discrete time steps $\{t_{j}\}_{j=0}^{N_{\mathrm{T}}}$. Define
@@
Quantization.
Each entry $u(x_{i},t_{j})$ is mapped to
$Q_{i,j}=\begin{cases}500,&\text{if }u_{\max}=u_{\min},\\ \displaystyle\mathrm{round}\Bigl(150+(u(x_{i},t_{j})-u_{\min})\,\frac{850-150}{u_{\max}-u_{\min}}\Bigr),&\text{otherwise},\end{cases}$
@@
Reconstruction.
To recover an approximate floating-point value $\tilde{u}(x_{i},t_{j})$ from $Q_{i,j}$, we apply a linear reconstruction map that approximates the original values:
$\tilde{u}(x_{i},t_{j})=\begin{cases}u_{\min},&\text{if }u_{\max}=u_{\min},\\ \displaystyle u_{\min}+(Q_{i,j}-150)\,\frac{u_{\max}-u_{\min}}{850-150},&\text{otherwise}.\end{cases}$
Quantization Error Floor.
We report two per-time-step error metrics for the reconstructed solution:
@@
LLM Inference Setup.
For LLM inference described in the Methods section of the main text, we consider only the token library $\mathcal{V}$ consisting of three-digit numbers (000–999) and the comma delimiter (,) that encodes spatial position. All other tokens outside $\mathcal{V}$ are masked, and sampling is renormalized over this set. This guarantees that every output corresponds to either a valid grid value entry or a spatial delimiter token, thereby preserving one-to-one alignment with the serialized PDE solution. For multi-step rollouts, temporal delimiters (semicolon) are inserted deterministically to mark time-step boundaries and appear in the LLM input context.
1.2 Spline-Based Random Initial Condition Construction
In the Results section of the main text, we construct each random initial condition $u_{0}(x)$ by sampling independent values on a fixed grid and then fitting a $C^{2}$ interpolant via cubic splines, yielding a function that is $C^{2}$ on [L,L][-L,L]. This analytic procedure is computationally cheap and helps ensure that any variation in LLM prediction error across different spatial discretizations arises purely from discretization, not from changes in the underlying random sample.
Coarse Grid and Sampling.
Let the 1D domain be [L,L][-L,L], with Dirichlet boundary values
@@ 1.2 Spline-Based Random Initial Condition Construction
Spline Interpolant.
Use SciPy’s CubicSpline with default “not-a-knot” end conditions to fit
@@
Resampling at a New Grid Resolution.
To evaluate different spatial discretizations, choose a new cardinality of interior points $N_{\mathrm{X}}^{\mathrm{new}}$ and set
@@
Parameter Choices.
All experiments in the Results section of the main text for the Allen–Cahn PDE are conducted with $N_{\mathrm{X}}=14,u_{\mathrm{BC}}=-1,a=-0.5,b=0.5$. Parameter choices for the additional PDEs are provided in Section 3.
2 Evaluation Metrics and Reference Solution Setup
@@
1:Fixed Quantities: Initial conditions $\{\tilde{u}_{0,m}\}_{m=1}^{M}$ and corresponding reference solutions $\{\tilde{u}_{m}(x,t)\}_{m=1}^{M}$ precomputed on a suitably highly refined finite-difference grid using appropriate schemes (FTCS for Allen–Cahn and Fisher–KPP; BTCS for heat; leapfrog for wave).
2:Monte Carlo Trials: For $m=1,2,\ldots,M$, run $\mathrm{A}\in\{\mathrm{LLM},\mathrm{Classical\penalty 10000\ Solver}\}$ given $\{\tilde{u}_{m}(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}-1}$ and obtain prediction $\{\hat{u}_{m}^{\mathrm{A}}(x_{i},t_{N_{\mathrm{T}}})\}_{i=1}^{N_{\mathrm{X}}}$.
@@
$\hskip-28.45274pt\bar{H}^{\mathrm{LLM}}_{m}=-\frac{1}{N_{\mathrm{X}}}\sum_{i=1}^{N_{\mathrm{X}}}\sum_{y\in\mathcal{V}}p(y\mid x_{i},N_{\mathrm{T}})\log p(y\mid x_{i},N_{\mathrm{T}}),\qquad\text{(Mean Entropy)}$
@@
In this section, we detail the Monte Carlo procedure used to compute the evaluation metrics in the Results section of the main text. For completeness, we also report results using the Maximum Absolute Error (MaxAE), which captures the worst-case deviation:
@@
As shown in Fig. S1, MaxAE exhibits qualitatively similar trends to the RMSE reported in the Results section of the main text. Algorithm 1 summarizes the Monte Carlo procedure used for computing evaluation metrics in the one-step prediction task. The same procedure extends naturally to the multi-step setting, where the metrics are computed identically at each predicted time slice.
@@
Figure S1: Prediction accuracy of Llama-3 models evaluated using the MaxAE metric, under the same experimental setup as in the Results section of the main text.
3 Results on Additional PDEs
3.1 Results on Fisher–KPP, Heat, and Wave Equations
We extend the analysis from the Results section of the main text to three additional PDEs: the Fisher–KPP equation, the heat equation, and the wave equation. All experiments (one-step prediction, multi-step rollout, and entropy-based uncertainty quantification) are performed under the same setup as in the Results section of the main text, with homogeneous Dirichlet boundary conditions unless noted otherwise. For the Fisher–KPP equation, we set $a=0.2,b=0.8$ (instead of a=0.5,b=0.5a=-0.5,b=0.5), following Section 1.2 so that the initial condition $u(x,0)$ lies within $[0,1]$, consistent with interpreting $u(x,t)$ as a population density. For the wave equation, which is second-order in time, we additionally impose zero initial velocity, i.e., $\partial_{t}u(x,0)=0$, so that the dynamics are fully determined by the initial condition $u(x,0)$. Explicitly, the governing equations are:
@@ 3.1 Results on Fisher–KPP, Heat, and Wave Equations
For the Fisher–KPP equation, we adopt commonly used parameter values with diffusion coefficient $D=0.002$ and reaction rate $r=1$ [42, 43]. In the representative results (Figs. S2S5), we present results with thermal diffusivity $k=0.01$ and wave speed $c=0.2$. In Fig. S6, we further confirm that the qualitative scaling trends persist across a range of $k$ and $c$ values. For each PDE, we additionally include representative numerical benchmarks to contextualize LLM predictions: FTCS and IMEX for the Fisher–KPP equation, FTCS and BTCS (backward time, centered space) for the heat equation, and leapfrog and Crank–Nicolson for the wave equation [35]. Beyond Dirichlet boundaries, we also study the heat equation under homogeneous Neumann boundary conditions, where total thermal energy conservation is the key structural property. Notably, LLM rollouts preserve this conservation law, suggesting that ICL can capture deeper invariants of PDE dynamics. Full details are provided in Subsection 3.2.
As shown in Figs. S2, S3, S4, and S5, the qualitative trends closely mirror those observed for the Allen–Cahn equation presented in the Results section of the main text. In the one-step prediction setting, accuracy improves systematically with longer temporal context while degrading at finer spatial discretizations. In the multi-step rollout setting, errors accumulate algebraically with the rollout horizon, resembling the global error growth of classical numerical solvers. Entropy-based analysis reveals a consistent three-stage progression in ICL behavior, and prediction uncertainty increases with longer spatial outputs.
@@
Figure S2: One-step prediction error for the Fisher–KPP, heat, and wave equations as a function of input context length, under the experimental setup of the Results section of the main text.
@@
Figure S3: One-step prediction error for the Fisher–KPP, heat, and wave equations as a function of output length, under the experimental setup of the Results section of the main text.
@@
Figure S4: Multi-step prediction error for the Fisher–KPP, heat, and wave equations as a function of prediction steps, under the experimental setup of the Results section of the main text.
@@
Figure S5: Uncertainty analysis of the Fisher–KPP, heat, and wave equations, under the experimental setup of the Results section of the main text.
@@
Figure S6: One-step prediction error for the Llama-3.2-3B model across varying PDE parameters, under the experimental setup of the Results section of the main text.
Figure S7: Relative energy deviation $\Delta\mathrm{E}$ over prediction steps for LLM rollouts compared against classical finite difference solvers (FTCS, BTCS) under homogeneous Neumann boundary conditions for the heat equation. Shaded regions denote 95% confidence intervals over 20 random initial conditions. The grid-induced reference is computed by evaluating the total thermal energy with a high-resolution solver and then restricting it to the same coarse spatial grid used in the experimental setup; this represents the minimal deviation expected from discretization alone. Llama-3.1-8B predictions remain close to this reference and exhibit substantially lower conservation error than coarse-grid finite difference solvers across the rollout horizon.
3.2 Conservation Properties in the Heat Equation with Neumann Boundaries
Motivation.
Beyond continuing spatiotemporal trajectories, an important question is whether LLMs internalize deeper invariants of PDE dynamics. The heat equation with homogeneous Neumann boundary conditions and no internal source term offers a natural test case: it models an insulated rod, where no heat can flow across the boundaries. In this setting, the total thermal energy is conserved for all time [20]. Remarkably, we find that LLM rollouts respect this conservation law more faithfully than coarse-grid finite difference solvers under the same setup, suggesting that ICL captures structural properties of the dynamics rather than performing naive extrapolation.
Conservation Law.
The governing PDE is:
@@ 3.2 Conservation Properties in the Heat Equation with Neumann Boundaries
Relative Energy Deviation.
To evaluate conservation in rollouts, we approximate $\mathrm{E}(t)$ using the trapezoidal rule on the spatial grid $\{x_{i}\}_{i=0}^{N_{\mathrm{X}}+1}$. The grid follows the uniform setup in the main section:
@@
Figure S7 reports results under the same setup as the Neumann-boundary heat equation experiments shown in Fig. S4. The only modification is that initial conditions are drawn from $a=0,b=1$ (instead of a=0.5,b=0.5a=-0.5,b=0.5), following Subsection 1.2, so that the conserved energy $\mathrm{E}(0)$ is bounded away from zero. This ensures numerical stability when evaluating relative deviations. We also verified that one-step, multi-step, and uncertainty-evolution analyses under Neumann boundaries with $a=0,b=1$ show consistent qualitative behavior with the results in Subsection 3.1 (which use a=0.5,b=0.5a=-0.5,b=0.5); to avoid redundancy, we do not reproduce these plots here, but the full results are available in the accompanying GitHub repository. Under this setup, LLM rollouts maintain thermal energy close to its conserved value across time, and notably more faithfully than coarse finite difference solvers operating at the same resolution.
@@
4 Results for Instruction-Tuned Llama-3 Variants
@@
Figure S8: Prediction accuracy of instruction-tuned Llama-3 models, using the same experimental setup described in the Results section of the main text.
@@
Figure S9: Uncertainty analysis of instruction-tuned Llama-3 models, using the same experimental setup described in the Results section of the main text. Top: Mean spatial entropy $\bar{H}$ as a function of context length $N_{\mathrm{T}}$ (left) and output length $N_{\mathrm{X}}$ (right). Bottom: Token-level softmax distributions at the syntax-only stage for Llama-3.1-8B-Instruct. The overall entropy behavior remains consistent with the pretrained base model, with one notable difference: at very short contexts ($N_{\textrm{T}}=2$), the instruction-tuned 8B variant exhibits higher average spatial entropy, reflecting greater uncertainty under minimal context. Separator tokens (e.g., commas) are still predicted with near-perfect confidence, and this increased uncertainty arises from spatial value tokens more frequently acting as generic placeholders rather than producing deterministically incorrect outputs, as observed in the base model.
@@
We replicate the full experimental setup from the Results section of the main text using instruction-tuned variants of Llama-3 models: Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, and Llama-3.2-1B-Instruct. These models are instruction-tuned for assistant-like chat, whereas the pretrained base models are designed to support a broader range of natural language generation tasks [16]. All evaluations (one-step prediction, multi-step rollout, and entropy-based uncertainty quantification) are conducted using the same setup as described in the Results section of the main text.
@@
As shown in Fig. S8, the qualitative trends closely mirror those of the base models. In particular, in one-step prediction settings, accuracy improves with longer input context and degrades with increasing output length. In multi-step prediction settings, predictions exhibit algebraic error accumulation. Similarly, Fig. S9 shows that prediction uncertainty undergoes stage-wise transitions as input context increases and grows with output length. Notably, under the same accuracy evaluation setup, the smaller instruction-tuned models, Llama-3.2-3B-Instruct and Llama-3.2-1B-Instruct, show reduced prediction accuracy compared to their pretrained base counterparts. This observation is consistent with prior findings from [9], which suggest that alignment procedures such as instruction tuning and Reinforcement Learning with Human Feedback (RLHF) can adversely affect time-series forecasting performance in Llama-2 models. In contrast, we do not observe such a negative impact on the 8B instruction-tuned variant, indicating that newer, larger models may be more robust to the effects of alignment in the context of continuing the spatiotemporal dynamics of PDEs.
5 Results for Other Model Families